# Large-scale Long-tailed Disease Diagnosis on Radiology Images

### v2

@@ 9 Conclusion
In this paper, we focus on solving the problem of multi-modal, multi-label long-tailed case-level diagnosis task. Specifically, we propose a new large-scale diagnosis dataset, namely, RP3D-DiagDS, with 39,026 cases (192,675 scans) labeled with detailed disorders covering 930 ICD-10-CM. On model design, we propose one unified architecture that supports both 2D and 3D input, together with a fusion module, to integrate information from multiple scans of one patient. Additionally, we adopt knowledge enhancement training strategy, leveraging the rich medical domain knowledge to improve the radiologic diagnosis performance. Our final train model also shows strong transferring ability to various external datasets regardless of their imaging modalities, shooting anatomies and target classes. We believe this work can serve as a transition phase between the specialist and generalist models, offering a more feasible and targeted playground for exploring sophisticated algorithms in academia labs, providing a unique opportunity for detailed error analysis, which is often impractical in the development of large-scale generalist models due to prohibitive computational costs.
10 Data and Code Availability
Our dataset RP3D-DiagDS can be found in https://huggingface.co/datasets/QiaoyuZheng/RP3D-DiagDS and our codes can be found in https://github.com/qiaoyu-zheng/RP3D-Diag

### v3

@@ 5 CONCLUSION
In this work, we mainly focus on the diagnosis of general diseases (rare diseases included) based on radiology images, which can be formulated as a problem of multi-modal, multi-label long-tailed case-level diagnosis task. Specifically, we exploit the online data resources, constructing a new large-scale diagnosis dataset, namely, RP3D-DiagDS, with 40,936 cases (195,010 scans) labeled with detailed disorders covering 930 ICD-10-CM. On model design, we propose one unified architecture that supports both 2D and 3D input, together with a fusion module, to integrate information from multiple scans of one patient. Additionally, we adopt knowledge enhancement training strategy, leveraging the rich medical domain knowledge to improve the radiological diagnosis performance. Our final model demonstrates strong performance on large-scale radiological disease diagnosis, and shows strong transferring ability to various external datasets regardless of their imaging modalities, shooting anatomies, and target classes. As a result, we show that publicly shared medical data on the Internet is a tremendous and valuable resource that can potentially support building a generalist model towards AI4healthcare.
6 DATA AVAILABILITY
Due to the Radiopaedia license  [^4] https://radiopaedia.org/licence?lang=us, we do not directly provide the original dataset. To obtain the data, readers must first get the permission from Radiopaedia’s founders, then we will share the download link. We also provide the model checkpoint in https://huggingface.co/QiaoyuZheng/RP3D-DiagModel following the apache-2.0 open-source license and all related result numbers can be reproduced based on the provided model checkpoint.
7 CODE AVAILABILITY
Our codes can be found in https://github.com/qiaoyu-zheng/RP3D-Diag under the MIT License, with all training, evaluation, fine-tuning, and zero-shot evaluation information.